19. When opening a door, you push with a force with magnitude F and which is directed perpendicularly to the door. If you apply the force closer to the hinges of the door.. A. the torque you exert and the angle the door moves through for a given distance which your B. the torque you exert will be larger, but the angle the door moves through for a given distance C. there will be no change in the angle the door moves through for a given distance which your D. the torque you exert and the angle the door moves through for a given distance which your E. the torque you exert will be smaller, but the ange the door moves through for a given hand moves will both be smaller which your hand moves will be smaller. hand moves. hand moves will both be larger distance which your hand moves will be largerWhy is the answer E and not A?
Answer #1

As torque is given as , T= F*d (*applied force is p[erpendicular)

F= force applied to open th door

d= distance

As the distance is small near the hinges, torque exterted is also small. so torque is small.

if the force applied is not perpendicular, T= F*d* sin \theta

For the given distance, d is constant, and to obtain same torque, angle  \theta should be larger.